# MTL-KD: Multi-Task Learning Via Knowledge Distillation for Generalizable Neural Vehicle Routing Solver

### v2

@@ Abstract
Multi-Task Learning (MTL) in Neural Combinatorial Optimization (NCO) is a promising approach to train a unified model capable of solving multiple Vehicle Routing Problem (VRP) variants. However, existing Reinforcement Learning (RL)-based multi-task methods can only train light decoder models on small-scale problems, exhibiting limited generalization ability when solving large-scale problems. To overcome this limitation, this work introduces a novel multi-task learning method driven by knowledge distillation (MTL-KD), which enables the efficient training of heavy decoder models with strong generalization ability. The proposed MTL-KD method transfers policy knowledge from multiple distinct RL-based single-task models to a single heavy decoder model, facilitating label-free training and effectively improving the model’s generalization ability across diverse tasks. In addition, we introduce a flexible inference strategy termed Random Reordering Re-Construction (R3C), which is specifically adapted for diverse VRP tasks and further boosts the performance of the multi-task model. Experimental results on 6 seen and 10 unseen VRP variants with up to 1000 nodes indicate that our proposed method consistently achieves superior performance on both uniform and real-world benchmarks, demonstrating robust generalization abilities.
@@ 1 Introduction
The Vehicle Routing Problem stands as a recognized classical challenge within the field of combinatorial optimization. Due to its widespread applicability across numerous real-world scenarios [1, 2, 3], including logistics distribution, traffic scheduling, and emergency response, the VRP has consistently been a subject of intense research interest. However, traditional optimization algorithms [4, 5, 6] often necessitate intricate domain-specific knowledge for their design and exhibit limitations in efficiency when applied to large-scale instances, thereby hindering their scalability in practical applications.
In recent years, Neural Solvers [7, 8, 9, 10] have demonstrated significant potential in tackling VRPs. These models offer high computational efficiency and reduced reliance on domain-specific knowledge, achieving competitive performance with traditional algorithms. However, the inherent design of these single-task models significantly restricts their versatility in handling diverse problem instances. The necessity to retrain the model even for slightly different problem settings severely limits their practical utility.
@@
The diverse nature of VRPs in the real world has spurred recent research efforts to develop unified models capable of solving multiple VRP variants and exhibiting good generalization on unseen tasks[11, 12, 13]. Currently, the majority of unified models for VRP variants adopt a heavy encoder-light decoder (HELD) architecture. These models have relatively low computational costs and can be directly trained using RL, demonstrating excellent performance on small-scale problems. However, their generalization ability significantly deteriorates when dealing with large-scale instances, primarily because the light decoder struggles to extract sufficient information from the high-density and complex node embeddings[14]. Furthermore, training them on large-scale instances also presents the challenge of excessively large gradient information, collectively leading to the inadequacy of existing VRP variant models in handling large-scale problems.
@@
On the other hand, the light encoder-heavy decoder[15, 16, 17] architecture has demonstrated excellent generalization capabilities on large-scale problems. Its iterative decoding process, where the heavy decoder continuously re-evaluates the relationships between the remaining nodes, coupled with the dynamic change in the number of nodes, contributes to the strong scale generalization performance of this architecture. To achieve scale generalization within a multi-task unified model, we aim to leverage the heavy decoder architecture. However, the substantial memory and computational demands inherent in heavy decoders render direct RL impractical; they typically employ supervised learning, which necessitates extensive labeled data, a demand further amplified by multi-task training. While SIT[17] employs a self-improved Training method using operations like local reconstruction [15] to refine model-generated solutions as training labels, this suffers from the generation of numerous low-quality labels during early and mid-training, leading to a protracted process and an increased training burden due to time-consuming pseudo-label generation. We provide a comprehensive literature review on autoregressive neural solvers and multi-task neural solvers for VRPs in Appendix A.
@@
To address the limitations of existing supervised training methods for heavy decoders, we introduce knowledge distillation as a novel training paradigm for the heavy decoder model, thereby circumventing the need for extensive label generation. We employ easily RL-trained HELD teacher models as supervision to train our multi-task heavy decoder model, thus achieving label-free training for the heavy decoder. The heavy decoder model trained with this method demonstrates strong generalization performance across both scales and tasks. Furthermore, during inference, we propose a general Random Reordering Reconstruction strategy for various VRP variants. By randomly reordering the external order of subtours, R3C significantly enhances solution sampling diversity, mitigates the risk of getting trapped in local optima, and further improves performance. Our contributions are summarized as follows: 1) We achieve efficient label-free training of heavy decoder models for multi-task VRPs through knowledge distillation. 2) We propose a novel R3C strategy to further enhance the performance of the multi-task model. 3) The MTL-KD model demonstrates excellent performance on seen training tasks, unseen tasks, and real-world datasets, exhibiting good scale generalization ability and significantly outperforming existing multi-task VRP models.
@@ 2 Preliminaries
By introducing additional constraints, CVRP can be extended into various variants. The four types of additional constraints studied in this paper include: (1) Open Route (O): Vehicles are not required to return to the depot after completing their service. (2) Time Windows (TW): Each node $i$ has a service time window $[s_{i},l_{i}]$, within which vehicles must arrive at the customer (early arrivals require waiting until the earliest service start time). (3) Backhaul (B): Customer demands can be positive (linehauls) or negative (backhauls), and there is no restriction on the sequence of visiting linehauls and backhauls customers. (4) Duration Limit (L): The total travel distance of each vehicle must not exceed a predefined threshold $L$. Combining these constraints with CVRP can form 16 different VRP variants, with specific details provided in Appendix B.
@@
Constructive neural solvers [7, 8, 15] typically employ an encoder-decoder architecture and construct solutions in an autoregressive manner. For a given VRP instance $\mathcal{G}=(V,E)$, where $V$ denotes the set of nodes, including node features such as demand, and $E$ represents the set of edges, often formed by a distance matrix. The encoder is used to extract node information and generate embedding vectors for each node, while the decoder constructs the next node based on the current partial solution state $s_{t-1}$. The decoder is subject to node constraints during the construction process to avoid generating infeasible solutions. The entire solving process can be represented as:
$\pi_{\theta}(\tau|\mathcal{G})=\prod_{t=1}^{\ell}\pi_{\theta}(a_{t}|s_{t},\mathcal{G}),$
where $\tau$ is the complete solution, $a_{t}$ is the action (node) selected at time step $t$, and $\ell$ is the total number of actions. The decoder continuously constructs nodes until all nodes have been visited, forming a complete solution. Through training, the model learns how to effectively select the next node based on the current partial solution, thereby optimizing the quality of the entire solution.
@@
Knowledge Distillation[18] is a model compression and knowledge transfer technique aimed at transferring the knowledge learned by a high-performing, pre-trained teacher model to a smaller student model. The core idea is to achieve this transfer by minimizing the discrepancy between the outputs of the teacher and student models. The student model’s training loss typically includes a distillation loss ($L_{KD}$) that measures the difference between the student’s predictions and the teacher’s soft targets, generally the KL divergence $L_{KD}=\text{KL}(\pi_{T}||\pi_{S})$. Additionally, depending on the specific task requirements, an original task loss ($L_{Task}$) can optionally be included, and the total loss ($L$) for the student model can be expressed as:
$L=\alpha L_{KD}+(1-\alpha)L_{Task}\quad$
where ($\alpha$) is a weight parameter. In the field of Neural Combinatorial Optimization, knowledge distillation has been used to achieve model lightweighting and can improve the generalization ability of models on unseen problems to a certain extent[19, 20, 21, 22].
@@ 3.2 Teacher Model Training
For each seen VRP variant task, we first generate the corresponding instance data. Then, we independently train a teacher model for each task, which adopts the POMO[8] structure and serves as a policy network. We optimize the parameters of the teacher models using the policy gradient method in RL, with the objective of maximizing the reward (negative tour length). The loss function[23] can be expressed as:
$L(\theta^{T})=-\mathbb{E}_{\tau\sim\pi_{\theta^{T}}(\cdot|\mathcal{G})}[(r(\tau)-b^{T}(\mathcal{G}))\log\pi_{\theta^{T}}(\tau|\mathcal{G})]$
where $b^{T}(\mathcal{G})$ is the baseline (average reward of multi-start trajectories for the instance), $\pi_{\theta^{T}}$ is the teacher model’s policy distribution, $\theta^{T}$ represents the parameters, and $r$ is the reward.
@@ 3.3 Student Model Training
During the student model training phase, we construct a multi-task heavy decoder model. In each training batch, we simultaneously generate instance data for the $N$ seen problem types and feed them into the student model for processing. At each decoding step $t$, the student model outputs a predicted probability distribution over the next nodes $\pi_{\theta^{S}}(a_{t}|s_{t},\mathcal{G})$, encompassing the probability distributions for the $N$ problem types. Concurrently, we input each problem instance to its corresponding pre-trained teacher model to obtain the teacher model’s probability distribution $\pi_{\theta^{T}n}(a_{t}|s_{t},\mathcal{G})$ under the same state $s_{t}$. We calculate the Kullback-Leibler divergence between the student model’s output probability distribution and the corresponding teacher model’s output probability distribution as the knowledge distillation loss $L_{KD}^{(t)}$:
$L_{KD}^{(t)}=\sum_{n=1}^{N}\text{KL}(\pi_{\theta^{T}n}(a_{t}|s_{t},\mathcal{G})\|\pi_{\theta^{S}}(a_{t}|s_{t},\mathcal{G}))$
By minimizing this loss function, the student model learns to mimic the behavior of multiple teacher models, thereby acquiring more generalized knowledge and improving its performance in multi-task learning and cross-task generalization. After each decoding step, we select the next node based on the student model’s probability distribution, transition to the new state, and repeat the aforementioned alignment process.
@@ 4.1 Encoder
Given an instance $S$ of the VRP, which includes $N$ node features $(s_{1}...s_{n})$, in VRP variants, the node features are uniformly represented as $(x,d,\delta,s,l)$, denoting the coordinate information, demand, service time, and the start and end times of the time window, respectively. These node features are then mapped through a linear layer to obtain the initial embedding $h^{0}=(h_{1}^{0}...h_{n}^{0})$. Subsequently, these initial embeddings are passed through a Transformer layer to capture node relationships and generate the node embeddings. Therefore, the output of the encoder can be represented as:
$H=\text{TF}(h^{0})=(h_{1},...,h_{n})$
where $H$ is the matrix of final node embeddings, and $h_{i}$ represents the embedding of the $i$-th node.
@@ 4.2 Decoder
During the decoding phase, we first extract the dynamic features $D=\{l,t,d,o\}$, encompassing the current vehicle load, current time, remaining route duration, and a binary flag indicating an open route. These dynamic features are combined with last visited node($h_{last}$) and the depot node($h_{depot}$) respectively, and then each combination is passed through its own linear layer. The resulting embeddings, along with the unvisited nodes($h_{unvisited}$), are processed by an $L$-layer Transformer network to yield updated node embeddings:
$h^{L}=\text{TF}([h_{unvisited},\text{Linear}(D\oplus h_{last}),\text{Linear}(D\oplus h_{depot})])$
Next, we compute compatibility scores between the embeddings of all unvisited nodes and a context embedding $h_{q}=\text{Linear}(\text{concat}(h_{last},h_{depot}))$ using single-head attention:
$c(h_{i}^{L},h_{q})=\text{SHA}(h_{i}^{L},h_{q})$
To ensure valid routes, a mask vector mask is applied to these scores. Finally, the probability $\pi(i|s_{t})$ of selecting the $i$-th unvisited node is determined by the softmax function applied to the masked compatibility scores:
$\pi(i|s_{t})=\text{softmax}(c(h_{i}^{L},h_{q})+\text{mask}_{i})$
where $c(h_{i}^{L},h_{q})$ is the compatibility score, $h_{i}^{L}$ is the embedding of the $i$-th unvisited node from $h^{L}$, and $s_{t}$ is the current decoder state. The mask vector contains 0 for valid unvisited nodes and $\infty$- ∞ for invalid ones.
To address the varying number of unvisited nodes across different instances within a batch, which can impact batch processing efficiency, we utilize padding to unify the sequence length and a masking mechanism to remove the influence of the padded depot node. Additionally, layer normalization [15] has been removed from all attention layers.
@@ 5 Random Reordering Re-Construct
The R3C strategy is detailed in Figure 3. Given a solution, we first decompose it into subtours, then randomize their external order in the sequence. Subsequently, a random-length contiguous segment is randomly sampled and re-optimized by our model. Improved segments replace the original. By randomizing subtour order, R3C allows for more random combinations of subtours into partial solutions, aiding escape from local optima and improving iterative performance. All sampled segments end at the depot. Additionally, feasible subtour reversals are also incorporated to further enhance search diversity.
@@ 6 Experiments
The baseline methods are categorized into two groups: traditional heuristic solvers and multi-task neural solvers. (1) Traditional Solvers: PyVRP [24]: An open-source Hybrid Genetic Search (HGS)-based solver supporting multiple complex VRP variants. It represents the current state-of-the-art in heuristic approaches for the problems considered in this study; OR-Tools [6]:Google’s open-source optimization toolkit, applicable to all VRP variants examined in this study. (2) Multi-task Neural Solvers: MT-POMO [11]: Multi-task extension of the POMO framework, enabling simultaneous learning across different VRP variants; MVMoE[12]: An architecture that enhances model capacity and improves performance over MT-POMO by incorporating Mixture-of-Experts (MoE) modules. Given the differences in problem variants and visible task training among the baseline models, we only compare methods with the same training setup.
@@
We adopte the same problem settings as MVMoE[12]. We generate a test set comprising 16 VRP problems across four scales: 100, 200, 500, and 1000. The dataset at the 100-scale contains 1,000 instances, while the others contain 128 instances each. In these settings, the vehicle capacity is uniformly set to 50, the duration limit is set to 3, the time window for the depot node in Time Window problems is [0, 4.6], and for backhaul problems, the proportion of backhaul nodes is set to 20%. Our datasets are generated based on the MVMoE codebase to evaluate the performance of various baselines.
@@
We follow a similar setup to MVMoE, training our model on 6 VRPs and evaluating its zero-shot generalization performance on 10 unseen tasks (refer to Figure 1). The training process consists of two phases: teacher model pre-training and student model training.(1) Teacher Model Pre-training: We employ the POMO model as our teacher model. For each of the 6 visible tasks, we independently traine a single-task model using reinforcement learning. All teacher models share the same parameter configuration: 6 encoder layers, 1 decoder layer, an embedding dimension of 128, 8 attention heads, and a Feedforward hidden dimension of 512. The teacher models are trained for 4000 epochs on random instances of scale 100, with 20,000 training instances per epoch, a batch size of 128, and an initial learning rate of 0.0001. (2) Student Model Training: The student model consists of 1 encoder layer and 6 decoder layers, with embedding dimensions set to 128 and 96, employing 8 multi-head attention heads and a Feedforward hidden dimension of 512. The student model simultaneously learns from the 6 pre-trained teacher models on datasets with a problem scale of 100. The training batch size is set to 1500 (250 * 6), with 24,000 training instances per epoch, an initial learning rate of 0.0001, and a total of 900 training epochs. After the 300th epoch, the learning rate is halved every 100 epochs.
@@
We test the results of a Single Trajectory Greedy Search (S.T.) on all methods. Furthermore, to compare with the results of other solvers exploring multiple trajectories under data augmentation (denoted as M×$\times$×aug8, where M is the number of nodes and aug8 represents 8 types of data augmentation[8]), our method employs the R3C strategy iterated 200 times under a single trajectory

### v3

@@ Abstract
Multi-Task Learning (MTL) in Neural Combinatorial Optimization (NCO) is a promising approach for training a unified model capable of solving multiple Vehicle Routing Problem (VRP) variants. However, existing Reinforcement Learning (RL)-based multi-task methods can only train light decoder models on small-scale problems, exhibiting limited generalization ability when solving large-scale problems. To overcome this limitation, this work introduces a novel multi-task learning method driven by knowledge distillation (MTL-KD), which enables efficient training of heavy decoder models with strong generalization ability. The proposed MTL-KD method transfers policy knowledge from multiple distinct RL-based single-task models to a single heavy decoder model, facilitating label-free training and effectively improving the model’s generalization ability across diverse tasks. In addition, we introduce a flexible inference strategy termed Random Reordering Re-Construction (R3C), which is specifically adapted for diverse VRP tasks and further boosts the performance of the multi-task model. Experimental results on 6 seen and 10 unseen VRP variants with up to 1,000 nodes indicate that our proposed method consistently achieves superior performance on both uniform and real-world benchmarks, demonstrating robust generalization abilities. The code is available at https://github.com/CIAM-Group/MTLKD.
@@ 1 Introduction
The Vehicle Routing Problem (VRP) is a classical combinatorial optimization problem (COP) with widespread application across numerous real-world scenarios, including logistics distribution, traffic scheduling, and emergency response [1, 2, 3, 4, 5]. Obtaining exact VRP solutions is computationally difficult due to its NP-hard nature [6]. Traditional heuristic algorithms [7, 8, 9] exhibit superior performance on many classic VRP tasks. However, they typically necessitate substantial domain-specific knowledge for design. In recent years, neural combinatorial optimization (NCO) [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] has emerged as a promising approach for tackling VRPs. This approach can automatically learn problem-solving strategies using neural networks and has the potential to minimize reliance on specialized domain knowledge. This field has developed rapidly, with some NCO algorithms approaching or even surpassing the performance of classical heuristic algorithms on some COPs [21, 22]. However, when solving different tasks, many NCO methods require modifying the model components for each individual task and then retraining the model, which significantly restricts their versatility in handling diverse VRP variants.
@@
Recently, considerable research efforts have been devoted to developing unified models capable of solving multiple VRP variants using Multi-Task Learning (MTL) [23, 24, 25]. The majority of unified models for VRP variants adopt a heavy encoder-light decoder (HELD) architecture. These models have relatively low computational costs and can be directly trained using RL, demonstrating excellent performance on small-scale problems. However, their generalization ability significantly deteriorates when solving with large-scale instances, primarily because the light decoder struggles to extract sufficient information from the high-density and complex node embeddings [26].
@@
On the other hand, the light encoder-heavy decoder [27, 28, 29] architecture demonstrates excellent generalization capabilities on large-scale problems. Its heavy decoder contributes to its strong scale generalization performance, which re-evaluates the relationships between the remaining nodes and first and last nodes during the iterative decoding process. Leveraging the heavy decoder architecture has the potential to achieve good scale generalization using a multi-task unified model. However, the substantial memory and computational demands inherent in heavy decoders render the model training using Reinforcement Learning (RL) impractical. Employing Supervised Learning (SL) is also difficult due to the absence of labeled data for multiple VRP variants. While SIT [29] employs a self-improved Training method using operations like local reconstruction [27] to refine model-generated solutions as training labels, it suffers from the generation of numerous low-quality labels during early and mid-training, leading to a protracted process and an increased training burden. We provide a comprehensive literature review on multi-task neural solvers for VRPs in Appendix A.
@@
To address the limitation associated with the training for the heavy decoder-based model on multiple VRP variants, we introduce a novel multi-task learning method driven by knowledge distillation (MTL-KD). The proposed MTL-KD method transfers policy knowledge from multiple distinct RL-based single-task models to a single heavy decoder model, facilitating label-free training and effectively improving the model’s generalization ability across diverse tasks. Furthermore, during inference, we propose a general Random Reordering Reconstruction strategy for various variants of the VRP. By randomly reordering the external order of subtours, R3C significantly enhances solution sampling diversity, mitigates the risk of getting trapped in local optima, and further improves performance. Our contributions are summarized as follows: 1) We achieve efficient label-free training of heavy decoder models for multi-task VRPs through the proposed MTL-KD method. 2) We propose a novel R3C strategy to further enhance the performance of the multi-task model. 3) The MTL-KD model demonstrates excellent performance on 6 seen training tasks and 10 unseen tasks with up to 1,000 nodes, and real-world datasets, exhibiting good scale generalization ability and significantly outperforming existing multi-task VRP models.
@@ 2 Preliminaries
By introducing additional constraints, CVRP can be extended into various variants. The four types of additional constraints studied in this paper include: (1) Open Route (O): Vehicles are not required to return to the depot after completing their service. (2) Time Windows (TW): Each node $i$ has a service time window $s_{i},l_{i}]$, within which vehicles must arrive at the customer (early arrivals require waiting until the earliest service start time). (3) Backhaul (B): Customer demands can be positive (linehauls) or negative (backhauls), and there is no restriction on the sequence of visiting linehauls and backhauls customers. (4) Duration Limit (L): The total travel distance of each vehicle must not exceed a predefined threshold $L$. Combining these constraints with CVRP can form 16 different VRP variants, with specific details provided in Appendix B.
@@
Constructive neural solvers [10, 11, 27] typically employ an encoder-decoder architecture and construct solutions in an autoregressive manner. For a given VRP instance $\mathcal{G}=(V,E)$, where $V$ denotes the set of nodes, including node features such as demand, and $E$ represents the set of edges, often formed by a distance matrix. The encoder is used to extract node information and generate embedding vectors for each node, while the decoder constructs the next node based on the current partial solution state $s_{t-1}$. The decoder is subject to node constraints during the construction process to avoid generating infeasible solutions. The entire solving process can be represented as: π_θ(τ|G)= ∏_t=1^ℓ π_θ(a_t|s_t,G), where $\tau$ is the complete solution, $a_{t}$ is the action (node) selected at time step $t$, and $\ell$ is the total number of actions. The decoder continuously constructs nodes until all nodes have been visited, forming a complete solution. Through training, the model learns how to effectively select the next node based on the current partial solution, thereby optimizing the quality of the entire solution.
@@
Knowledge Distillation [30] is a model compression and knowledge transfer technique aimed at transferring the knowledge learned by a high-performing, pre-trained teacher model to a smaller student model. The core idea is to achieve this transfer by minimizing the discrepancy between the outputs of the teacher and student models. The student model’s training loss typically includes a distillation loss ($L_{KD}$) that measures the difference between the student’s predictions and the teacher’s soft targets, generally the KL divergence $L_{KD}=\text{KL}(\pi_{T}||\pi_{S})$. Additionally, depending on the specific task requirements, an original task loss ($L_{Task}$) can optionally be included, and the total loss ($L$) for the student model can be expressed as: L = αL_KD + (1 - α) L_Task,   where ($\alpha$) is a weight parameter. In the field of NCO, knowledge distillation has been used to achieve model lightweighting and can improve the generalization ability of models on unseen problems to a certain extent [31, 32, 33, 34].
@@ 3.2 Teacher Model Training
For each seen VRP variant task, we first generate the corresponding instance data. Then, we independently train a teacher model for each task, which adopts the POMO [11] structure and serves as a policy network. We optimize the parameters of the teacher models using the policy gradient method in RL, with the objective of maximizing the reward (negative tour length). The loss function [35] can be expressed as: L(θ^T) = - E_τπ_θ^T(⋅|G) [(r(τ) - b^T(G)) logπ_θ^T(τ|G)], where $b^{T}(\mathcal{G})$ is the baseline (average reward of multi-start trajectories for the instance), $\pi_{\theta^{T}}$ is the teacher model’s policy distribution, $\theta^{T}$ represents the parameters, and $r$ is the reward.
@@ 3.3 Student Model Training
During the training phase of the student model, we employ a multi-task heavy decoder architecture that processes instances from all $N$ seen problem types simultaneously. For each training batch, we generate instances from all $N$ seen problem types and process them together. At each decoding step $t$, the student model outputs a probability distribution $\pi_{\theta^{S}}(a_{t}|s_{t},\mathcal{G})$ over the next node to be selected.
For knowledge distillation, we feed each problem instance to its corresponding pre-trained teacher model to obtain the teacher’s output distribution $\pi_{\theta^{T}n}(a_{t}|s_{t},\mathcal{G})$ at the same decoding step, ensuring that each problem instance is supervised by its corresponding specialized teacher model.
The knowledge distillation loss is computed as the Kullback-Leibler divergence between the student model’s output distribution and the corresponding teacher model’s output distribution: L_KD^(t) = ∑_n=1^N KL(π_θ^Tn(a_t|s_t,G) ∥ π_θ^S(a_t|s_t,G)).
By minimizing this loss function, the student model learns to mimic the behavior of multiple teacher models simultaneously, thereby acquiring generalized knowledge from different problem types. After each decoding step, the algorithm select the next node based on the student model’s probability distribution, transition to the new state, and repeat this knowledge distillation process.
@@ 4.1 Encoder
Given an instance $S$ of the VRP, which includes $N$ node features $(s_{1}...s_{n})$, in VRP variants, the node features are uniformly represented as $(x,d,\delta,s,l)$, denoting the coordinate information, demand, service time, and the start and end times of the time window, respectively. These node features are then mapped through a linear layer to obtain the initial embedding $h^{0}=(h_{1}^{0}...h_{n}^{0})$. Subsequently, these initial embeddings are passed through a Transformer layer to capture node relationships and generate the node embeddings. Therefore, the output of the encoder can be represented as: H = TF(h^0) = (h_1, …, h_n), where $H$ is the matrix of final node embeddings, and $h_{i}$ represents the embedding of the $i$-th node.
@@ 4.2 Decoder
During the decoding phase, we first extract the dynamic features $D=\{l,t,d,o\}$, encompassing the current vehicle load, current time, remaining route duration, and a binary flag indicating whether the route is open. These dynamic features are combined with the last visited node($h_{last}$) and the depot node($h_{depot}$) respectively, and then each combination is passed through its own linear layer. The resulting embeddings, along with the unvisited nodes($h_{unvisited}$), are processed by an $L$-layer Transformer network to yield updated node embeddings: h^L = TF([h_unvisited, Linear(D ⊕h_last), Linear(D ⊕h_depot)]). Next, we compute compatibility scores between the embeddings of all unvisited nodes and a context embedding $h_{q}=\text{Linear}(\text{concat}(h_{last},h_{depot}))$ using single-head attention: c(h_i^L, h_q) = SHA(h_i^L, h_q). To ensure valid routes, a mask vector is applied to these scores. Finally, the probability $\pi(i|s_{t})$ of selecting the $i$-th unvisited node is determined by the softmax function applied to the masked compatibility scores: π(i|s_t) = softmax(c(h_i^L, h_q) + mask_i), where $c(h_{i}^{L},h_{q})$ is the compatibility score, $h_{i}^{L}$ is the embedding of the $i$-th unvisited node from $h^{L}$, and $s_{t}$ is the current decoder state. The mask vector contains 0 for valid unvisited nodes and -$\infty$ for invalid ones.
To address the varying number of unvisited nodes across different instances within a batch, which can impact batch processing efficiency, we utilize padding to unify the sequence length and a masking mechanism to remove the influence of the padded depot node. Additionally, layer normalization [27] is removed from all attention layers.
@@ 5 Random Reordering Re-Construct
The R3C strategy is detailed in Figure 3. Given a solution, we first decompose it into subtours, then randomize their external order in the sequence. Subsequently, a random-length contiguous segment is randomly sampled and re-optimized by our model. If the re-optimized segment improves the objective value, it replaces the original segment. By randomizing subtour order, R3C allows for more random combinations of subtours into partial solutions, aiding escape from local optima and improving iterative performance. All sampled segments end at the depot. Additionally, feasible subtour reversals are also incorporated to further enhance search diversity.
@@ 6 Experiments
The baseline methods are categorized into two groups: traditional heuristic solvers and multi-task neural solvers. (1) Traditional Solvers: PyVRP [36] is an open-source Hybrid Genetic Search (HGS)-based solver supporting multiple complex VRP variants. It represents the current state-of-the-art in heuristic approaches for the problems considered in this study. OR-Tools [9] is Google’s open-source optimization toolkit, applicable to all VRP variants examined in this study.
(2) Multi-task Neural Solvers: MT-POMO [23] is a multi-task extension of the POMO framework, enabling simultaneous learning across different VRP variants. MVMoE [24] is an architecture that enhances model capacity and improves performance over MT-POMO by incorporating Mixture-of-Experts (MoE) modules. RouteFinder [25] is a multi-task training framework that enhances training efficiency through mixed batch training and enables fast adaptation to unseen tasks via lightweight adapter layers. The framework includes three model variants: RF-Transformer, RF-POMO, and RF-MVMoE. CaDA [37] incorporates constraint awareness and a dual-branch structure to better handle diverse tasks within a multi-task learning framework.
@@
We adopt the same problem settings as MVMoE [24]. We generate a test set comprising 16 VRP problems across four scales: 100, 200, 500, and 1000. The dataset at the 100-scale contains 1,000 instances, while the others contain 128 instances each. In these settings, the vehicle capacity is uniformly set to 50, the duration limit is set to 3, the time window for the depot node in Time Window problems is [0, 4.6], and for backhaul problems, the proportion of backhaul nodes is set to 20%. Our datasets are generated based on the MVMoE codebase to evaluate the performance of various baselines.
@@
We follow a similar setup to MVMoE, training our model on 6 VRPs and evaluating its zero-shot generalization performance on 10 unseen tasks (refer to Figure 1). The training process consists of two phases: teacher model pre-training and student model training.(1) Teacher Model Pre-training: We employ the POMO model as our teacher model. For each of the 6 visible tasks, we independently train a single-task model using reinforcement learning. All teacher models share the same parameter configuration: 6 encoder layers, 1 decoder layer, an embedding dimension of 128, 8 attention heads, and a Feedforward hidden dimension of 512. The teacher models are trained for 4000 epochs on random instances of scale 100, with 20,000 training instances per epoch, a batch size of 128, and an initial learning rate of 0.0001. (2) Student Model Training: The student model consists of 1 encoder layer and 6 decoder layers, with embedding dimensions set to 128 and 96, employing 8 multi-head attention heads and a Feedforward hidden dimension of 512. The student model simultaneously learns from the 6 pre-trained teacher models on datasets with a problem scale of 100. The training batch size is set to 1500 (250 * 6), with 24,000 training instances per epoch, an initial learning rate of 0.0001, and a total of 850 training epochs. After the 300th epoch, the learning rate is halved every 100 epochs.
@@
[…]